Suppose the manager of a construction supply house determined from historical records that the lead time demand for sand averages 49 tons. In addition, suppose the manager determined that demand during lead time could be described by a normal distribution that has a mean of 49 tons and a standard deviation of 6 tons. If the manager is willing to accept a service risk of no more than 3 percent:

What value of z is appropriate?
How much safety stock should be held?
What reorder point should be used?
Answer #1

Answer:


a) Lead time demand = 49

Standard deviation = 6

Acceptable stock-out risk = 3%

Hence, service level = 100 - 3% = 97%

1.88 (the value of z is given, by using the normal table, 97% fill-rate = 1.88 z-score)

b) Safety Stock is given by:

Safety Stock = Standard Deviation x Z

Therefore:

Safety Stock = 6 x 1.88

= 11.28 tons

c) ROP = Lead time demand + Safety Stock

Therefore:

ROP = 49 + 11.28

= 60.28 tons


answered by: Subrahmanyam golla